OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x + 8y = 16 -2x + 5y = 55

OpenStudy (anonymous):

Lets eliminate x by multiplying the second equation by 2 4x + 8y = 16 -4x +10y = 110 and adding like terms 18y = 126 What is the next step?